OpenStudy (swissgirl):

Let p be a prime natural number. Prove that \( (\mathbb{Z_p} -\{0\}, \cdot) \)is an associative algebraic system with identity 1 Use Eulicid's Lemma

OpenStudy (swissgirl):

i have no clue what Eulicid's lemma is

OpenStudy (kinggeorge):

Merely that if \(p\) is prime, and \(p|ab\), then \(p|a\) or \(p|b\).

OpenStudy (kinggeorge):

So we just need to show that it's associative here correct?

OpenStudy (swissgirl):

yup

OpenStudy (kinggeorge):

I just realized I wasn't doing it a good way. Let me think about this a little bit.

OpenStudy (swissgirl):

okkkkkkkkkkkkkkkkkkkkkkkkkkkkkkk

OpenStudy (kinggeorge):

Maybe this will work better. Let \(s\) be prime. Suppose \(s|a(bc)\). Then we have two cases. \(s|a\) or \(s|bc\). We can split up the second case into two additional cases; \(s|b\) or \(s|c\). Case 1: \(s|a\). Then \(s|ab\), so \(s|(ab)c\). Case 2: \(s|b\). Then \(s|ab\), so \(s|(ab)c\). Case 3: \(s|c\). Then \(s|(ab)c\). Using a similar argument, we can see that \(s|a(bc)\) iff \(s|(ab)c\). Therefore, these two numbers have the same unique factorization, so they are indeed the same number mod p. This shows associativity.

OpenStudy (swissgirl):

This seems perfect to meee

OpenStudy (kinggeorge):

Sorry, tab just crashed. The next step is to prove the identity is 1. This follows from associativity and the fact that 1 is the identity within ordinary multiplication in integers.

OpenStudy (kinggeorge):

\[1\cdot(a\cdot(b\cdot c))=(1\cdot a)\cdot(b\cdot c)=a\cdot(b\cdot c)\]

OpenStudy (swissgirl):

yaaa that is clear

OpenStudy (swissgirl):

Thanks @KingGeorge

OpenStudy (kinggeorge):

You're welcome.

OpenStudy (swissgirl):

dw i wont bombard you with questions

OpenStudy (kinggeorge):

I probably won't notice anyways. I've got to go again, and who knows when I'll be back on. I wasn't kidding when I said I'll be busy in the next couple weeks.

OpenStudy (swissgirl):

lol sounds like it is stressfull, Good Luck

OpenStudy (kinggeorge):

You have no idea :| I was busy from 8AM-11PM today. Tomorrow it'll start at 9, and hopefully end a little earlier. Anyways, see you later.

OpenStudy (swissgirl):

See ya around